OpenStudy (anonymous):

6) Rewrite with positive exponents. Assume that even roots of non negative quantities and that all denominators are nonzero. (2x/y^2)^-4

OpenStudy (anonymous):

ans = (y^2/2x) ^ 4 using x^a = 1/ x^-a